### v3

@@ 5 High-Scalability Quantum Computers
5.1 Super-fast Quantum Machines: The “quantum supremacy” of quantum machines over conventional computers proves that the former can do very computationally intensive jobs on a conventional computer far more quickly. In the quantum world, “quantum advantage” is an additional important phrase. A more realistic concept would be “quantum advantage”, which deals with solving a practical, real-world issue that cannot be effectively addressed on a traditional computer, as opposed to the theoretical “quantum supremacy” that would imply resolving a challenging issue on any conventional processor [2]. Quantum superiority has been shown, but finding real-world problems that quantum computers can effectively tackle remains an unexplored field of study mainly due to the decoherence of quantum bits. Most of the current generation of quantum computers is cumbersome and underpowered due to the materials used, which must be maintained at superconducting (extremely low) temperatures; yet, the promise of prospective commercial quantum computers is undeniable [52]. The current popularity of traditional computers and their meteoric rise in the 1950s provide the impetus for the possible advantages of industrial quantum computers. Older classical computers were cumbersome and required constant cooling, just like modern quantum computers. We may theoretically expect strong commercial quantum systems to attain “quantum advantage” in the not-too-distant future, much as the Artificial Intelligence (AI) concept began to take shape during the initial stages of traditional computing devices, even though these machines couldn’t have possibly handled the computations needed for AI [53].
@@
5.4 Scalable Commercial Quantum Computing Manufacturing Challenges: The biggest technical problem that needs to be solved before an industrial-grade quantum machine can be fully functional is noise or decoherence, which makes quantum processing mistakes (destroys entanglement of qubits) and stops quantum benefits from working. Until a qubit can be utilised, its starting state must be established, and gates and networks must also be developed. Even though photons maintain their coherent state for an extended period of time, it is difficult to construct quantum circuits using them. Companies like IBM, Google, Rigetti, and others are building quantum machines using quantum circuits based on superconductivity. Unfortunately, there is still a need to develop strategies for error correction or moderation due to the poor fidelity of these qubits, especially in two-qubit operations. If your quantum circuit utilises five or fewer qubits, you may build and operate it on IBM’s five-qubit cloud processor, which was made publicly available in 2016. In addition to their newly revealed 433-qubit quantum computer, IBM now provides cloud usage of quantum machines with up to 65 qubits.
5.5 Presently Accessible Infrastructure: In 2016, IBM unveiled its five-qubit IBM Quantum Experience quantum computer [57]. Along with the system’s release, a user manual and an interactive chat were made available. Rights to engage via quantum assembly language, a user-friendly interface, and a simulation extension were among the many features introduced to the IBM Quantum Experience later in 2017 [58]. After that, IBM released Qiskit, a tool that enhanced quantum processor coding. In addition, they established the quantum awards program and created a system with sixteen qubits. Superconducting qubits housed in a dilution refrigerator constitute the hardware of IBM’s quantum computers. The quantum composer is the name of the application’s user interface (GUI) that consumers engage with. When writing quantum assembly code, quantum composer is the tool of choice. Quantum experiments and algorithms may be more easily developed with the help of the GUI. One can also choose to use a simulator instead of a real Quantum Processing Unit (QPU). To run quantum computations through their paces, Rigetti Computing provides a Forest framework as a cloud-based quantum computing utility. A quantum processor from Forest has over 36 qubits, and it is possible to utilize Python to do hybridized conventional and quantum computations. The European cloud computing provider QuTech offers the quantum platform Quantum Inspire as part of its service offering. Without investing in or constructing a physical quantum computer, users can access the processing power of quantum algorithms using cloud-based quantum computing platforms.
@@ 6 Widening the Debate: New Trends and Potential Challenges
In light of the current study, we have been able to pinpoint a number of topics in quantum computing that are still being studied. Simulating complicated quantum processes has been the focus of much study, and post-quantum cryptography is now at its pinnacle. Fig. 2 summarises the main findings and recommendations that can be utilised by future researchers to further quantum computing research. In the realm of quantum technology, new fields of study are taking shape, including automation, handling energy, computer security, decentralised quantum computing, complicated mathematical chemistry and drug design [2]. It could take over a decade for these domains to fully implement quantum computing when they are first introduced. People have unrealistically high hopes for isothermal quantum computing, quantum management, and quantum security. Assuming they fall within the ambit of quantum computing, their development is anticipated to take short time [59]. There has been an excess of optimism around several areas of quantum technology, including the Internet, error-corrected quantum technology, digital information exploration, quantum-aided AI, and quantum-based satellite communications [60]. We have uncovered several unanswered questions and potential avenues for further study, all of which are subject to ongoing investigation on a worldwide scale.
@@
6.1 Technological and Developmental Challenges: The primary problem with quantum technology is its vulnerability, which arises from two main sources: 1) the fact that some qubits have a very short coherence period (which is very qubit technology dependent) since, due to their superconductivity, they lose their data extremely often. 2) Developing a quantum computer with minimal errors is challenging since quantum processes are unreliable because of the relatively substantial rate of errors needing a huge number of qubits for error handling. Additionally, error correction in quantum technology is far more difficult than in conventional computing due to the following reasons: (a) quantum errors are ongoing (including the two magnitudes and stages), (b) it is not possible to replicate unknown quantum states, and (c) evaluation may degrade a quantum state and erase the information in qubits. A large number of physical qubits are needed to execute a quantum algorithm successfully; this necessitates a tight and constant link between the classical structure and the quantum device, which in turn creates a massive control burden. Additionally, the connection and overhead costs increase the complexity of the run-time control, design, and installation for quantum computing processes. At the moment, the qubit count serves as a measure of quantum computing equipment’s computational capacity. However, this metric is off by a significant margin, and it raises questions about the viability of supercomputer-level quantum machines with over a thousand qubits. Qubit design necessitates an efficient cooling component to manage heat, which AI-driven systems may be able to do. This increases scalability and allows for the solution of dynamically scaled, tricky issues.
@@
6.2 Resilient and Sustainable Quantum Modelling: Since the actual application of quantum error mitigation remains a matter of wide debate, it is difficult to achieve trustworthy and fault-tolerant quantum computers. The sensitive nature of quantum states necessitates operating bits at extremely cold temperatures and requires highly precise manufacture [43]. Accurately measuring the full quantum state is similarly difficult, making verification a difficult task. When compared to conventional computing, the likelihood of calculation mistakes is much higher. Quantum structures cannot function properly without a reliable method of error correction. In order to facilitate better verification of exact manufacturing restrictions, further reevaluation of quantum communication infrastructure is required. However, due to strict tolerances and the need to prevent using poorly positioned qubits to minimise error, testing qubits after manufacture is a challenging task. To achieve sufficient reliability to enable sustained quantum computation, iterative error mitigation is required [36]. To provide trustworthy service in the years to come, state-of-the-art AI/ML-based methods may be utilised for automatic error identification and rectification on the fly [61]. Nonetheless, it results in additional expenses for training AI/ML methods [62].
@@
6.7 Simulations for Advanced Quantum Research: In the near future, small-scale ”quantum simulators” with 50–100 qubits of computing power may be accessible, allowing quantum computers to model complicated biological, physical, and chemical issues [56]. To comprehend and utilise quantum technology, it is necessary to combine the knowledge of several experts with the essentials of conventional computing [53]. In addition, quantum simulators can mimic the natural system and solve complicated issues in a controlled environment, allowing researchers to study the interplay of several parameters—questions that would be impossible to accomplish using conventional or supercomputer systems. When developing quantum computers, simulators can make use of entanglement and superposition, two of their key features [58]. To conduct large-sized and complicated operations connected to biology and chemistry with optimum outcomes, the scalability of simulations needs to be increased in the future.
@@
Summary of Findings and Take-Aways: There are several unanswered questions and some good ideas for where to go from here. To date, it has been unclear how to combine these performance features into a single quantum computing approach. In order to construct a quantum computer capable of concurrent activities, a quantum computing approach that enables quantum I/O to have all the required classified properties is important. A post-quantum cryptography system is developed to safeguard conventional cryptographic basics and protocols by using the computational power of a quantum computer, which can solve mathematical issues in milliseconds. In order to make symmetrical cryptography basics and algorithms more resistant to the widely-known quantum assaults, post-quantum cryptography was developed. Additionally, the difficulties in scaling up the number of qubits that have been actually realised thus far mean that modern commercial quantum computers have yet to be capable of replacing conventional supercomputers. It is uncertain when that may occur. There is currently no clear indication of when quantum computers will begin to supplant conventional computers in difficult tasks, despite the fact that the next decade will be absolutely thrilling for industrial quantum computing. Even if quantum computing does become feasible, digital supercomputers will continue to exist as a complement to potential quantum computers. The question of how to effectively operate an algorithm with quantum properties is a critical one for designers. There is significant control overhead due to the high number of physical qubits that are necessary, which in turn require constant and tight communication between the classical substrate and the quantum device. Due to the ongoing issue of the correction of quantum errors, it is difficult to accomplish trustworthy and resilient quantum calculations. The sensitive nature of quantum states necessitates operating bits at extremely cold temperatures and precise manufacture. Additionally, using quantum computing to manage a massive dataset with an extensive number of gadgets (100–1000 qubits) might enhance the effectiveness and scalability of AI methods. To realistically apply hybrid computing (quantum and conventional computing) and tackle today’s most difficult business challenges, further effort is required. The advent of quantum computing will have far-reaching benefits for many other areas, including computer security, biology, economics, and the production of new substances.
7 Conclusions
@@
This article offers a vision and identifies various potential challenges on the topic of quantum computing. It has been found that entanglement and superposition, two quantum mechanics instances, are anticipated to be crucial for resolving computer issues. We discussed a number of quantum software methods and technologies, industrial quantum computers, and cryptography after quantum computers. Finally, we highlight a number of concerns that have yet to be solved, as well as promising new prospects for research and development in the field of quantum technology.

### v4

@@ 5 High-Scalability Quantum Computers
5.1 Super-fast Quantum Machines: The “quantum supremacy” of quantum machines over conventional computers proves that the former can do very computationally intensive jobs on a conventional computer far more quickly. In the quantum world, “quantum advantage” is an additional important phrase. A more realistic concept would be “quantum advantage”, which deals with solving a practical, real-world issue that cannot be effectively addressed on a traditional computer, as opposed to the theoretical “quantum supremacy” that would imply resolving a challenging issue on any conventional processor [2]. Quantum superiority has been shown, but finding real-world problems that quantum computers can effectively tackle remains unsolved mainly due to the decoherence of quantum bits. Most of the current generation of quantum computers is cumbersome and underpowered due to the materials used, which must be maintained at superconducting (extremely low) temperatures; yet, the promise of prospective commercial quantum computers is undeniable [68]. The current popularity of traditional computers and their meteoric rise in the 1950s provide the impetus for the possible advantages of industrial quantum computers. Older classical computers were cumbersome and required constant cooling, just like modern quantum computers. We may theoretically expect strong commercial quantum systems to attain “quantum advantage” in the not-too-distant future, much as the Artificial Intelligence (AI) concept began to take shape during the initial stages of traditional computing devices, even though these machines couldn’t have possibly handled the computations needed for AI [69].
@@
5.4 Scalable Commercial Quantum Computing Manufacturing Challenges: The biggest technical problem that needs to be solved before an industrial-grade quantum machine can be fully functional is noise or decoherence, which makes quantum processing mistakes (destroys entanglement of qubits) and stops quantum computing benefits. Until a stable qubit can be realised, its starting state must be established, and gates and networks must also be developed. Even though photons maintain their coherent state for an extended period of time, it is difficult to construct quantum circuits using them. Companies like IBM, Google, Rigetti, and others are building quantum machines using quantum circuits based on superconductivity. Unfortunately, there is still a need to develop strategies for error correction or moderation due to the poor fidelity of these qubits, especially in two-qubit operations. If a quantum circuit utilises five or fewer qubits, we can build and operate it on IBM’s five-qubit cloud processor, which was made publicly available in 2016. In addition to their newly revealed 433-qubit quantum computer, IBM now provides cloud usage of quantum machines with up to 65 qubits.
5.5 Presently Accessible Infrastructure: In 2016, IBM unveiled its five-qubit IBM Quantum Experience quantum computer [73]. Along with the system’s release, a user manual and an interactive chat were made available. Rights to engage via quantum assembly language, a user-friendly interface, and a simulation extension were among the many features introduced to the IBM Quantum Experience later in 2017 [74]. After that, IBM released Qiskit, a tool that enhanced quantum processor coding. In addition, they established the quantum awards program and created a system with sixteen qubits. Superconducting qubits housed in a dilution refrigerator constitute the hardware of IBM’s quantum computers. The quantum composer is the name of the application’s user interface (GUI) that consumers engage with. When writing quantum assembly code, quantum composer is the tool of choice. Quantum experiments and algorithms may be more easily developed with the help of the Graphical User Interface (GUI). One can also choose to use a simulator instead of a real Quantum Processing Unit (QPU). To run quantum computations through their paces, Rigetti Computing provides a Forest framework as a cloud-based quantum computing utility. A quantum processor from Forest has over 36 qubits, and it is possible to utilize Python to do hybridized conventional and quantum computations. The European cloud computing provider QuTech offers the quantum platform Quantum Inspire as part of its service offering. Without investing in or constructing a physical quantum computer, users can access the processing power of quantum algorithms using cloud-based quantum computing platforms.
@@ 6 Widening the Debate: New Trends and Potential Challenges
In light of the current study, we have been able to pinpoint a number of topics in quantum computing that are still being investigated. Simulating complicated quantum processes has been the focus of much study, and post-quantum cryptography is now at its pinnacle. Fig. 2 summarises the main findings and recommendations that can be utilised by future researchers to further quantum computing research. In the realm of quantum technology, new fields of study are taking shape, including automation, handling energy, computer security, decentralised quantum computing, complicated mathematical chemistry and drug design [2]. It could take over a decade for these domains to fully implement quantum computing when they are first introduced. People have unrealistically high hopes for isothermal quantum computing, quantum management, and quantum security. Assuming they fall within the ambit of quantum computing, their development is anticipated to take short time [75]. There has been an excess of optimism around several areas of quantum technology, including the Internet, error-corrected quantum technology, digital information exploration, quantum-aided AI, and quantum-based satellite communications [76]. We have uncovered several unanswered questions and potential avenues for further study, all of which are subject to ongoing investigation on a worldwide scale.
@@
6.1 Technological and Developmental Challenges: The primary problem with quantum technology is its vulnerability, which arises from two main sources: 1) the fact that qubits have a very short coherence period (which is very qubit technology dependent) since, due to their superconductivity, they lose their data extremely often. 2) Developing a quantum computer with minimal errors is challenging since quantum processes are unreliable because of the relatively substantial rate of errors needing a huge number of qubits for error handling. Additionally, error correction in quantum technology is far more challenging than in conventional computing due to the following reasons: (a) quantum errors are ongoing (including the two magnitudes and stages), (b) it is not possible to replicate unknown quantum states, and (c) evaluation may degrade a quantum state and erase the information in qubits. A large number of physical qubits are needed to execute a quantum algorithm successfully; this necessitates a tight and constant link between the classical structure and the quantum device, which in turn creates a massive control burden. Additionally, the connection and overhead costs increase the complexity of the run-time control, design, and installation for quantum computing processes. At the moment, the qubit count serves as a measure of quantum computing equipment’s computational capacity. However, this metric is off by a significant margin, and it raises questions about the viability of supercomputer-level quantum machines with over a thousand qubits. Qubit design necessitates an efficient cooling component to manage heat, which AI-driven systems may be able to do. This increases scalability and allows for the solution of dynamically scaled, tricky issues.
@@
6.2 Resilient and Sustainable Quantum Modelling: Since the actual application of quantum error mitigation remains a matter of wide debate, it is difficult to achieve trustworthy and fault-tolerant quantum computers. The sensitive nature of quantum states necessitates operating bits at extremely cold temperatures and requires high precision manufacturing [44]. Accurately measuring the full quantum state is similarly difficult, making verification a difficult task. When compared to conventional computing, the likelihood of calculation mistakes is much higher. Quantum structures cannot function properly without a reliable method of error correction. In order to facilitate better verification of exact manufacturing restrictions, further reevaluation of quantum communication infrastructure is required. However, due to strict tolerances and the need to prevent using poorly positioned qubits to minimise error, testing qubits after manufacture is a challenging task. To achieve sufficient reliability to enable sustained quantum computation, iterative error mitigation is required [36]. To provide trustworthy service in the years to come, state-of-the-art AI/ML-based methods may be utilised for automatic error identification and rectification on the fly [77]. Nonetheless, it results in additional expenses for training AI/ML methods [78].
@@
6.7 Simulations for Advanced Quantum Research: In the near future, small-scale "quantum simulators" with 50–100 qubits of computing power may be accessible, allowing quantum computers to model complicated biological, physical, and chemical issues [72]. To comprehend and utilise quantum technology, it is necessary to combine the knowledge of several experts with the essentials of conventional computing [69]. In addition, quantum simulators can mimic the natural system and solve complicated issues in a controlled environment, allowing researchers to study the interplay of several parameters—questions that would be impossible to accomplish using conventional or supercomputer systems. When developing quantum computers, simulators can make use of entanglement and superposition, two of their key features [74]. To conduct large-sized and complicated operations connected to biology and chemistry with optimum outcomes, the scalability of simulations needs to be increased in the future.
@@ 7 Summary of Findings, Take-Aways and Conclusions
There are several unanswered questions and some good ideas for where to go from here. To date, it has been unclear how to combine these performance features into a single quantum computing approach. In order to construct a quantum computer capable of concurrent activities, a quantum computing approach that enables quantum I/O to have all the required classified properties is important. A post-quantum cryptography system is developed to safeguard conventional cryptographic basics and protocols by using the computational power of a quantum computer, which can solve mathematical issues in milliseconds. In order to make symmetrical cryptography basics and algorithms more resistant to the widely-known quantum assaults, post-quantum cryptography was developed. Additionally, the difficulties in scaling up the number of qubits that have been actually realised thus far mean that modern commercial quantum computers have yet to be capable of replacing conventional supercomputers. It is uncertain when that may occur. There is currently no clear indication of when quantum computers will begin to supplant conventional computers in difficult tasks, despite the fact that the next decade will be absolutely thrilling for industrial quantum computing. Even if quantum computing does become feasible, digital supercomputers will continue to exist as a complement to potential quantum computers. The question of how to effectively operate an algorithm with quantum properties is a critical one for designers. There is significant control overhead due to the high number of physical qubits that are necessary, which in turn require constant and tight communication between the classical substrate and the quantum device. Due to the ongoing issue of the correction of quantum errors, it is difficult to accomplish trustworthy and resilient quantum calculations. The sensitive nature of quantum states necessitates operating bits at extremely cold temperatures and precise manufacture. Additionally, using quantum computing to manage a massive dataset with an extensive number of gadgets (100–1000 qubits) might enhance the effectiveness and scalability of AI methods. To realistically apply hybrid computing (quantum and conventional computing) and tackle today’s most difficult business challenges, further effort is required. The advent of quantum computing will have far-reaching benefits for many other areas, including computer security, biology, economics, and the production of new substances.
@@
Finally, this article offers a vision and identifies various potential challenges on the topic of quantum computing. It has been found that entanglement and superposition, two quantum mechanics instances, are anticipated to be crucial for resolving computer issues. We discussed a number of quantum software methods and technologies, industrial quantum computers, and cryptography after quantum computers. Finally, we highlight a number of concerns that have yet to be solved, as well as promising new prospects for research and development in the field of quantum technology.